OpenStudy (twizttiez):

Which ordered pair is the solution to the system of equations? x+y=4 x-y=-6 A. (2, 2) B. (1, −7) C. (−1, 5) D. (−2, 4)

OpenStudy (twizttiez):

@Algorithmic

OpenStudy (anonymous):

x+y=4 x-y=-6 Add the equations 2x=-2 Solve for x x=-1 x+y=4 Substitute x=1 into one of the -1+y=4 original equations and solve for y y=1+4 y=5 x+y=4 Be sure to check your answer -1+5=4 in both equations 4=4 x-y=-6 -1-5=-6 -6=-6 The Solution is (-1,5)